OpenStudy (anonymous):

simplify: 3(to the 5th) times 3 times 3(to the 2nd)

OpenStudy (anonymous):

\[3^5 *3*3^2\] is it that?

OpenStudy (anonymous):

Yeah like that

OpenStudy (anonymous):

3 to the fifth is the same thing as 3*3*3*3*3. Basically, the to the 'x' part is asking how many 3's are multiplied together. In your case, you have 5 of them (3 to the 5th) then another one (times 3) and then 2 more (3 to the 2nd)

OpenStudy (anonymous):

The 'times 3' part is the same thing as 3 to the first power so it's the same idea.

OpenStudy (anonymous):

going off what he said. You basically have 3^8 since x^a*x^b=x^(a+b). In your case 3=x and a and b are the various powers you have which are 5,1,2. Your answer is 6561 in case you need to know.

OpenStudy (anonymous):

The options for answers are: 3 the the 8th 3 to the 10th 3 to the 7th 3 to the 9th

OpenStudy (anonymous):

3 to the 8th